# SAKURA: On the Multi-hop Reasoning of Large Audio-Language Models Based on Speech and Audio Information

### v2

@@ Abstract
Large audio-language models (LALMs) extend the large language models with multimodal understanding in speech, audio, etc. While their performances on speech and audio-processing tasks are extensively studied, their reasoning abilities remain underexplored. Particularly, their multi-hop reasoning, the ability to recall and integrate multiple facts, lacks systematic evaluation. Existing benchmarks focus on general speech and audio-processing tasks, conversational abilities, and fairness but overlook this aspect. To bridge this gap, we introduce SAKURA, a benchmark assessing LALMs’ multi-hop reasoning based on speech and audio information. Results show that LALMs struggle to integrate speech/audio representations for multi-hop reasoning, even when they extract the relevant information correctly, highlighting a fundamental challenge in multimodal reasoning. Our findings expose a critical limitation in LALMs, offering insights and resources for future research.
@@
Large language models (LLMs)[1, 2] have revolutionized AI research, extending beyond natural language processing to domain-specific applications like computer vision[3] and speech processing [4, 5]. This shift drives the rise of large multimodal models, e.g., large vision-language models (LVLMs)[6, 7] and large audio-language models (LALMs)[8, 9, 10, 11, 12, 13, 7], equipping LLMs with cross-modal understanding.††*Equal Contribution
Research has also explored LLMs’ fundamental reasoning abilities [14, 15, 16], with multi-hop reasoning, the skill of recalling and connecting multiple facts, recognized as a key to complex reasoning [17, 18, 19, 20]. Particularly, latent multi-hop reasoning [19, 20], which involves multi-hop reasoning over internal parameterized knowledge stored in LLMs, is crucial for efficiently utilizing learned knowledge without heavily relying on external retrieval. Multi-hop reasoning datasets also provide insights into LLMs’ internal mechanism [21], underscoring the importance of studying this ability.
Despite growing interest in reasoning, multi-hop reasoning of large audio-language models (LALMs) based on speech and audio remains unexplored. Existing benchmarks assess general speech and audio-processing performance [22, 23], conversational abilities [24, 25], and fairness/bias [26, 27], while studies on LALMs’ fundamental abilities [28, 29, 30] remain limited and do not specifically address multi-hop reasoning.
To address this gap, we introduce SAKURA (Speech and Audio-based Question-answering Benchmark for Multi-hop Reasoning of Large Audio-Language Models), a novel benchmark for systematic evaluation of LALMs’ multi-hop reasoning using speech and audio information. SAKURA has four tracks based on speech/audio attributes including speaker gender, spoken language, speaker emotion, and animal sound. Each track includes single-hop and multi-hop sub-tracks, both comprising 500 multiple-choice questions. Single-hop questions require direct perception of an attribute (e.g., “What is the animal in the sound?”), whereas multi-hop questions demand further reasoning based on the extracted information (e.g., “Which of the listed features aligns best with the physical appearance of the animal in the sound?”). This design compels models to recall and integrate the extracted auditory cues when tackling multi-hop questions. Figure 1 provides examples of single-hop and multi-hop reasoning paths. In total, SAKURA offers 4000 human-verified multiple-choice questions, providing not only a structured framework for evaluating LALMs’ speech/audio-based multi-hop reasoning but also a valuable resource for future research on their reasoning mechanisms.
@@ 2.1 Overview
SAKURA evaluates LALMs’ multi-hop reasoning with speech/audio information across four tracks, covering fundamental attributes: speaker gender (Gender), spoken language (Language), speaker emotion (Emotion), and animal sounds (Animal), chosen for their significance in speech/audio processing. Each track includes single-hop and multi-hop sub-tracks. Here, we define the single-hop reasoning of LALMs as the direct perception and extraction of speech/audio attributes, forming the foundation for speech/audio-based multi-hop reasoning.
@@ 2.3 Textual question-answer pair generation
Figure 1 illustrates our question-answer data generation pipeline. We used GPT-4o444gpt-4o-2024-11-20, https://openai.com to generate question-answer pairs. In the single-hop sub-tracks, questions directly queried target attributes. To enhance diversity, we used GPT-4o to paraphrase manually crafted questions, such as “What is the animal in the sound?”, producing multiple variations with differing phrasings and lengths. These paraphrased questions were then paired with speech/audio data, with the associated attribute label as the golden answer and distractors selected from alternative labels.
The generation process for the multi-hop sub-tracks was more complex. For each track, we used GPT-4o to generate several “tested topics” and three sample questions per topic to clarify their meaning. A topic is a concept that multi-hop questions can target based on the track’s attribute labels (e.g., “physical traits” in the Animal track). To ensure quality and diversity, we manually filtered the topics using these criteria:
@@
Relevance and Groundedness: Topics must be relevant to the track’s attribute and supported by the corresponding information. For example, topics like “determining the age of the speaker” in the Gender track or “determining the religious beliefs of the speaker” in the Language track were excluded.
@@
Objectivity: Topics must be objective and unbiased. For example, “the association of animals and human personality” was excluded due to cultural ambiguity.
@@
Uniqueness: Topics should be distinct. For example, only one of “feeding habits” and “food sources” could be included.
For each topic, GPT-4o generated diverse questions along with golden answer candidates and distractors for all attribute labels, forming a choice pool for pairing with speech/audio data. For example, for a question on “physical traits” with “cat” as the attribute label, GPT-4o might generate “claws” as a golden answer candidate and “tentacles” and “hooves” as distractors.
@@
Before finalizing the instructions, the choice order, including the golden answer and distractors, was shuffled to ensure the golden answer appeared equally frequent across positions (e.g., (a), (b), etc.), mitigating potential positional biases in LALMs (i.e., favoring a certain option, such as “(a)”, irrespective of content). The instruction was constructed by concatenating the question with the shuffled choices, with option prefixes (e.g., “(a)”, “(b)”) inserted. The paired speech/audio inputs, instructions, and golden answers were included in SAKURA.
@@ 3.1 Evaluation metrics
The final judgment should be summarized as a binary “correct/incorrect” label to facilitate post-processing.
@@ 3.2 Baselines
Cascaded systems comprised three components: a speech recognition (ASR) module, an audio captioning (AAC) module, and a text-based LLM, forming two variants: (1) “ASR+LLM,” where the LLM answered questions based only on ASR transcriptions, and (2) “ASR+AAC+LLM,” where audio captions enriched the input to the LLM. Specifically, we used Whisper-large-v3 [38] for ASR, Qwen2-Audio-Instruct for AAC, and LLaMA-3.1-8B-Instruct [1] as the LLM.
@@ 5 Conclusion, limitations, and future work
We introduce SAKURA, the first benchmark for systematically evaluating LALMs’ multi-hop reasoning with speech and audio. Our findings show that LALMs struggle to recognize certain speech and audio attributes, exhibiting perception blind spots. Even when accurately extracting relevant information, they fall short in speech/audio-based multi-hop reasoning despite good text-based reasoning skills, highlighting the need for better integration of multimodal information into their reasoning process.

### v3

@@ Abstract
Large audio-language models (LALMs) extend the large language models with multimodal understanding in speech, audio, etc. While their performances on speech and audio-processing tasks are extensively studied, their reasoning abilities remain underexplored. Particularly, their multi-hop reasoning, the ability to recall and integrate multiple facts, lacks systematic evaluation. Existing benchmarks focus on general speech and audio-processing tasks, conversational abilities, and fairness, but overlook this aspect. To bridge this gap, we introduce SAKURA, a benchmark assessing LALMs’ multi-hop reasoning based on speech and audio information. Results show that LALMs struggle to integrate speech/audio representations for multi-hop reasoning, even when they extract the relevant information correctly, highlighting a fundamental challenge in multimodal reasoning. Our findings expose a critical limitation in LALMs, offering insights and resources for future research.
@@
Large language models (LLMs)[1, 2] have revolutionized AI research, extending beyond natural language processing to domain-specific applications like computer vision[3] and speech processing [4, 5]. This shift drives the rise of large multimodal models, e.g., large vision-language models (LVLMs)[6, 7] and large audio-language models (LALMs)[8, 9, 10, 11, 12, 13, 7], equipping LLMs with cross-modal understanding.
Research has also explored LLMs' fundamental reasoning abilities [14, 15, 16], with multi-hop reasoning, the skill of recalling and connecting multiple facts, recognized as a key to complex reasoning [17, 18, 19, 20]. Particularly, latent multi-hop reasoning [19, 20], which involves multi-hop reasoning over internal parameterized knowledge stored in LLMs, is crucial for efficiently utilizing learned knowledge without heavily relying on external retrieval. Multi-hop reasoning datasets also provide insights into LLMs' internal mechanism [21], underscoring the importance of studying this ability.
Despite growing interest in reasoning, multi-hop reasoning of large audio-language models (LALMs) based on speech and audio remains unexplored. Existing benchmarks assess general speech and audio-processing performance [22, 23], conversational abilities [24, 25], and fairness/bias [26, 27], while studies on LALMs' fundamental abilities [28, 29, 30] remain limited and do not specifically address multi-hop reasoning.
To address this gap, we introduce SAKURA (Speech and Audio-based Question-answering Benchmark for Multi-hop Reasoning of Large Audio-Language Models), a novel benchmark for systematic evaluation of LALMs' multi-hop reasoning using speech and audio information. SAKURA has four tracks based on speech/audio attributes including speaker gender, spoken language, speaker emotion, and animal sound. Each track includes single-hop and multi-hop sub-tracks, both comprising 500 multiple-choice questions. Single-hop questions require direct perception of an attribute (e.g., ``What is the animal in the sound?"), whereas multi-hop questions demand further reasoning based on the extracted information (e.g., ``Which of the listed features aligns best with the physical appearance of the animal in the sound?"). This design compels models to recall and integrate the extracted auditory cues when tackling multi-hop questions. Figure 1 provides examples of single-hop and multi-hop reasoning paths. In total, SAKURA offers 4000 human-verified multiple-choice questions, providing not only a structured framework for evaluating LALMs’ speech/audio-based multi-hop reasoning but also a valuable resource for future research on their reasoning mechanisms.
@@ 2.1 Overview
SAKURA evaluates LALMs' multi-hop reasoning with speech/audio information across four tracks, covering fundamental attributes: speaker gender (Gender), spoken language (Language), speaker emotion (Emotion), and animal sounds (Animal), chosen for their significance in speech/audio processing. Each track includes single-hop and multi-hop sub-tracks. Here, we define the single-hop reasoning of LALMs as the direct perception and extraction of speech/audio attributes, forming the foundation for speech/audio-based multi-hop reasoning.
@@ 2.3 Textual question-answer pair generation
Figure 1 illustrates our question-answer data generation pipeline. We used GPT-4o444gpt-4o-2024-11-20, https://openai.com to generate question-answer pairs. In the single-hop sub-tracks, questions directly queried target attributes. To enhance diversity, we used GPT-4o to paraphrase manually crafted questions, such as ``What is the animal in the sound?", producing multiple variations with differing phrasings and lengths. These paraphrased questions were then paired with speech/audio data, with the associated attribute label as the golden answer and distractors selected from alternative labels.
The generation process for the multi-hop sub-tracks was more complex. For each track, we used GPT-4o to generate several ``tested topics" and three sample questions per topic to clarify their meaning. A topic is a concept that multi-hop questions can target based on the track's attribute labels (e.g., ``physical traits" in the Animal track). To ensure quality and diversity, we manually filtered the topics using these criteria:
@@
Relevance and Groundedness: Topics must be relevant to the track's attribute and supported by the corresponding information. For example, topics like ``determining the age of the speaker" in the Gender track or ``determining the religious beliefs of the speaker" in the Language track were excluded.
@@
Objectivity: Topics must be objective and unbiased. For example, ``the association of animals and human personality" was excluded due to cultural ambiguity.
@@
Uniqueness: Topics should be distinct. For example, only one of ``feeding habits" and ``food sources" could be included.
For each topic, GPT-4o generated diverse questions along with golden answer candidates and distractors for all attribute labels, forming a choice pool for pairing with speech/audio data. For example, for a question on ``physical traits" with ``cat" as the attribute label, GPT-4o might generate ``claws" as a golden answer candidate and ``tentacles" and ``hooves" as distractors.
@@
Before finalizing the instructions, the choice order, including the golden answer and distractors, was shuffled to ensure the golden answer appeared equally frequent across positions (e.g., (a), (b), etc.), mitigating potential positional biases in LALMs (i.e., favoring a certain option, such as ``(a)", irrespective of content). The instruction was constructed by concatenating the question with the shuffled choices, with option prefixes (e.g., ``(a)", ``(b)") inserted. The paired speech/audio inputs, instructions, and golden answers were included in SAKURA.
@@ 3.1 Evaluation metrics
The final judgment should be summarized as a binary ``correct/incorrect" label to facilitate post-processing.
@@ 3.2 Baselines
Cascaded systems comprised three components: a speech recognition (ASR) module, an audio captioning (AAC) module, and a text-based LLM, forming two variants: (1) ``ASR+LLM," where the LLM answered questions based only on ASR transcriptions, and (2) ``ASR+AAC+LLM," where audio captions enriched the input to the LLM. Specifically, we used Whisper-large-v3 [38] for ASR, Qwen2-Audio-Instruct for AAC, and LLaMA-3.1-8B-Instruct [1] as the LLM.
@@ 5 Conclusion, limitations, and future work
We introduce SAKURA, the first benchmark for systematically evaluating LALMs' multi-hop reasoning with speech and audio. Our findings show that LALMs struggle to recognize certain speech and audio attributes, exhibiting perception blind spots. Even when accurately extracting relevant information, they fall short in speech/audio-based multi-hop reasoning despite good text-based reasoning skills, highlighting the need for better integration of multimodal information into their reasoning process.
@@
6 Acknowledgement
We acknowledge the computational and storage support provided by the National Center for High-performance Computing (NCHC) of the National Applied Research Laboratories (NARLabs) in Taiwan.